The activation energy for the isomerization of cyc
Answer #1

Answer 7.

Expression for rate constant of given reaction (Arrhenius equation) can be written as follows:

k = Ae^{(-Ea/RT)}

Where, k = Rate constant

A = Arrhenius factor

Ea = activation energy = 274 kJ/mol

R = Universal gas constant = 0.008314 kJ/(mol.K)

T = temperature in Kelvin

Thus, expression for rate constant at two temperatures can be written as follows:

k_{1} = Ae^{(-Ea/RT_{1})}

k_{2} = Ae^{(-Ea/RT_{2})}

Where, k1 is the rate constant at temperature T1 and k2 is the rate constant at temperature T2

T1 = 257oC = (257 + 273.15) K = 530.15 K

T2= 280oC = (280 + 273.15) K = 553.15 K

Thus, the factor by which rate of reaction increases i.e. ratio of K2/K1 can be written as follows:

\frac{k_{2}}{k_{1}} = e^{(-Ea/RT_{2})+(Ea/RT_{1})}

\frac{k_{2}}{k_{1}} = e^{(-274/(0.008314)(553.15))+(274/(0.008314)(530.15))}

\frac{k_{2}}{k_{1}} = e^{(-59.58)+(62.16)}

\frac{k_{2}}{k_{1}} = e^{(2.58)}

\frac{k_{2}}{k_{1}} = 13.2

Thus, the factor by which rate of reaction increases is 13.2

--------------------------------------------------------------------------------------------------------------------------------------------------

Answer 8.

The approximate energy level diagram for catalyzed and uncatalyzed reaction can be shown as follows:

It can be seen from above diagram, the activation energy for uncatalyzed forward reaction is 101kJ/mol. The uncatalyzed reverse reaction will be endothermic reaction and its activation energy will be 198 kJ/mol as shown in above diagram.